OpenStudy (anonymous):

The temperature was 65 degrees at daybreak. Then it dropped 2 degrees per hour until dusk. If x is the number of hours since daybreak, this equation represents the temperature during the day: y = -2x + 65 OR y = 65 - 2x. Use the equation to predict the temperature 6 hours later

OpenStudy (agreene):

plug in 6 for x and you have your answer, -2(6)+65 = 53

OpenStudy (anonymous):

haha, those two equations are equivalent

OpenStudy (anonymous):

-2x + 65 = 65 - 2x